# 2]Dept. of Physics and Astronomy and LENS, Florence Univ., via Sansone 1, I-50019 Sesto Fiorentino, Italy

### v1

@@ 3.2 Noise robustness
Due to limited accessibility to real quantum hardware, computation of kernels is simulated without noise on a classical hardware, using tools from Qiskit library. To show how noise affects our results, we implement some models introducing depolarizing, bit-flip and amplitude damping errors [28].
First we consider the quantum SVC defined by the 6-qubits PQC. We select a splitting of the dataset in training and test set and we run the method without any noise. Then we add noise through the depolarizing error model, initially with error rate $p=0.01$, and after that we progressively increase $p$ by 0.01. For each run we compute the achieved accuracy. We repeat it also for bit-flip error and amplitude damping error, and subsequently we perform the same analysis for the 8-qubits and the 12-qubits methods. For the latter methods, excessive computational cost makes it necessary to reduce the number of samples before running the methods. In particular we consider 20 samples per label for the 8-qubits method, and 15 samples per label for the 12-qubits method.
In Fig. 6 we show how the accuracies of the 6-qubits SVC vary as the noise rate $p$ increases in the 3 different noise models. The accuracy curves stop at the $p$ value where the methods collapse, namely at $p^{*}$ such that for all $p\in[p^{*},1]$ the accuracy is the same as the one obtained at $p^{*}$. In the 8-qubits and 12-qubits SVCs, since the amount of noise injected grows with the number of qubits, the collapse happens for smaller $p^{*}$. All methods seems to be more robust to amplitude damping error than to depolarizing and bit-flip errors.
Figure 6: Accuracies (in percentage) of 6-qubits SVC as noise rate increases in the depolarizing, bit-flip and amplitude damping noise models.
@@ 4 Conclusion
The aim of this work is comparing the performances of classical and quantum kernel methods for classifying samples of AD patients and healthy people subjected to handwriting tests. To achieve this, we have used first some classical approaches and then we have compared their performances with the ones achieved by quantum methods defined by various PQCs. Amongst classical SVCs, the one defined by the radial basis function kernel achieves average accuracy of 86.76%, while linear kernel and sigmoid kernel obtain 84.46% and 85.14% respectively. The quantum SVC defined by the 6-qubits PQC gets to 85.95%, the one defined by the 8-qubits PQC reaches 88.51% and the one defined by the 12-qubits circuit achieves average accuracy of 87.70%. The fact that the latter does not outperform the model defined by the 8-qubits circuit seems to indicate that performance does not automatically improve by increasing the number of qubits, and some ansatze are better suited for smaller PQCs, hence more feasible with the currently available quantum processors.
@@
In addition to running the methods considering all features together we have also defined methods based on data subsampling. We have considered separately the features associated with the 3 categories of tasks, namely features derived from graphic tasks, copy tasks or memory and dictation tasks. With graphic features, RBF and 8-qubits SVC achieve average accuracies of 79.19% and 79.86% respectively. When considering copy features, both RBF and 8-qubits SVC achieve average accuracy of 83.11%. Using the features derived from the tasks in the category ‘memory and dictation’, RBF and 8-qubits SVC achieve average accuracies of 73.65% and 71.89% respectively. Then we have examined the features obtained from each of the 25 task individually, defining a model for each task, and we have used the majority vote decision rule to combine the predictions made by the 25 models. Classical and quantum SVC obtain mean accuracies of 80.41% and 81.49% respectively. When restricted to only the best 5 tasks, i.e. the tasks that have achieved singularly the best accuracies in the previous runs, both classical and quantum SVC obtain average accuracies of 87.16%.
Kernels in quantum methods have been calculated simulating quantum computation without noise. We have also investigated how different types of noise affect the performance of quantum methods. From our analysis it appears that the methods show good robustness to the error introduced by amplitude damping, which is associated to energy dissipation. In fact, quantum SVCs tend to perform poorly only for high value of the noise rate related to the damping probability. Both depolarization and bit flip error instead cause the methods to perform worse for smaller values of noise rate. However, on real quantum computers different sources of noise act simultaneously, due to interaction with the environment and interaction between qubits, or imperfect application of quantum gates [29]. Therefore quantum noise models, without in-depth physical experiments, can only partially tackle the complexity of this problem.
Our study shows that quantum methods can achieve better performances in classification tasks even involving big datasets, as for the case of medical data analysis. Nowadays only QPUs with a limited number of noisy qubits are available. However this does not seem to affect too much the capability of quantum kernel methods. Moreover the constant improvement of quantum technologies suggests that there is a lot of room for enhancements to make quantum methods achieve even better results.
The techniques used in our work may have a wide range of applications in QML algorithms for classification and regression, and can be implemented to enhance classical ML models for similar problems of early screening, e.g. detection of autism in children [30], or identification of various types of cancer [31, 32].

### v2

@@ 3.2 Noise robustness
Due to limited accessibility to real quantum hardware, computation of kernels is simulated without noise on a classical hardware. Now we compare the performance of the methods executed on a noiseless Aer Simulator with those run on a fake backend. A fake backend is a simulated quantum computing environment that mimics the behavior of real IBM Quantum systems, providing real hardware qubit properties to simulate realistic conditions. We select a splitting of the dataset in training and test set, leaving 20% of the samples for the latter, and we run the quantum SVCs on the Aer Simulator without noise, with number of shots equal to 256. This means that each circuit run is executed 256 times and the measurement results are aggregated to provide a more accurate representation of the quantum state. The default number of shots is usually set to 1024, but we choose to reduce this to speed up the computation, especially when using the fake backend. Then, with the same splitting and the same number of shots, we run the methods on Fake Melbourne backend, which simulates a 14-qubits quantum processor. The average T1 and T2 times of Fake Melbourne are 50 and 70 microseconds respectively, where T1 is a measure of how long a qubit can retain its state before it decays to ground state, and T2 indicates how long a qubit can stay in a superposition state before losing coherence, due to interactions with the environment [37]. A noiseless backend always yields the same result, whereas the performance of a fake backend, influenced by noise, is variable. Therefore we run the methods on the fake backend 20 times, storing their predictions and using the majority rule to yield the final prediction. In Fig. 5 we plot, for each of the 3 methods, the accuracies achieved in the noiseless run and the 20 noisy runs.
Figure 5: Accuracies of the quantum SVCs run on a noiseless simulator and on a noisy fake backend.
In Table 4 we compare the performance of the noiseless models and the noisy models combined with the majority vote decision rule.
Table 4: Accuracy (expressed in percentage) of quantum SVCs run on Aer Simulator without noise and on the noisy backend Fake Melbourne; for the latter the accuracy is obtained combining the predictions made on 20 executions with the majority vote decision rule.
6q 8q 12q
Aer Simulator 72.86 62.86 77.14
Fake Melbourne 77.14 64.29 77.14
The comparable or slightly better accuracy of the Fake Melbourne backend indicates that running multiple executions and using majority voting is a good strategy to mitigate variability caused by noise, achieving more consistent and accurate outcomes.
@@ 4 Conclusion
The aim of this work is comparing the performances of classical methods and quantum kernel methods for classifying samples of AD patients and healthy people subjected to handwriting tests. The classical models used are Support Vector Classification, k-Nearest Neighbors and Decision Tree; quantum SVCs are defined by ansatze of 6, 8 and 12 qubits. Among the classical methods, the SVC delivers the highest performance, albeit with significant inconsistency: average accuracy is 85.28% and standard deviation 7.46%. Both kNN and DT are significantly less accurate. Evaluating the quantum methods we find that 6 and 8-qubits SVCs perform similarly to the classical SVC with slightly lower accuracy but better consistency. The quantum SVC with 12 qubits emerges as the most effective model, demonstrating superior accuracy and consistency. This indicates that increasing the number of qubits can significantly boost the performance of quantum classifiers, potentially unlocking more advanced capabilities and better results.
@@
In addition to running the methods considering all features together we have also defined methods based on data subsampling. We have considered separately the features associated with the 3 categories of tasks, namely features derived from graphic tasks, copy tasks or memory and dictation tasks. In this case we just compare the performances of the 2 best models, i.e. classical SVC and 12-qubits SVC. The quantum algorithm generally shows improved or comparable performance across different feature subsets, highlighting its robustness and potential advantages in handling diverse types of data. Then we have examined the features obtained from each of the 25 task individually, defining a model for each task, and we have used the majority vote decision rule to combine the predictions made by the 25 models; after that we have considered just the best 5 tasks, i.e. the tasks that have achieved singularly the best accuracies in the previous runs. The quantum SVC demonstrates an advantage in accuracy over the classical SVC, whether considering all tasks or just the best ones, though it tends to have higher variability in its predictions.
Kernels in quantum methods have been calculated simulating quantum computation without noise. We have also investigated how noise affects the performance of quantum methods using a fake backend, which mimics the behavior of a real noisy backend. Averaging the predictions from several executions on the noisy backend, with the majority rule, leads to a slightly superior accuracy, counteracting the variability introduced by noise. Several studies in classical machine learning have explored how introducing randomness and noise can enhance the performance of kernel methods [38][39]. This concept holds significant promise when applied to quantum devices, which inherently operate in noisy environments. Noise may facilitate a broader exploration of the solution space, potentially leading to better solutions that might be missed in a noise-free environment. Regarding the use of real quantum hardware, currently, only IBM Quantum systems with 127 or more qubits are accessible, and each run for our computations takes 2 to 3 seconds. Similar considerations can be made for other quantum device providers. Despite computing multiple kernels in a single run, the execution of the methods exceeds the few minutes allocated in the free plan. Additionally, the number of kernels to compute increases quadratically with the number of samples. This makes quantum kernel methods more prone to time constraints compared to other quantum classifiers that can assign a label to a sample in a single run. Exploring optimizations, such as more efficient kernel computation techniques or hybrid quantum-classical approaches, could help mitigate these time constraints and make quantum kernel methods more feasible for real-world applications.
While our experimental results are competitive with the current state of the art, there remains potential for further enhancement. In the quantum circuits we have used, only one hyperparameter, the bandwidth, is tuned in the validation phase. Enhancing the results could be achieved by incorporating one or more layers of parametric gates with tunable hyperparameters for optimization. Moreover, parameters ‘C’ , ‘gamma’ and ‘tol’ could be also added in the parameter grid of quantum methods. To enhance the user-friendliness of quantum methods, particularly in the clinical diagnostic process, we aim to automate the implementation of the methods we have developed. By simplifying the process of choice of the ansatz, optimization of parameters and execution on fake or real backends, we wish to make quantum kernel methods more accessible and usable for a wide range of applications, improving the performance of classical ML models for similar problems of early screening, e.g. detection of autism in children [40], or identification of various types of cancer [41, 42].